Skip to content
CalcGospel 國際數學圖譜
返回

CIE 9709 2024 November Paper 51 Q1

A Level / CIE / S1

CIE 9709 2024 Nov Paper 51 Paper · Question 1

题目

Problem

Nicola throws an ordinary fair six-sided dice. The random variable XX is the number of throws that she takes to obtain a 6.

(a) Find P(X<8)\mathrm{P}(X<8).

(2)

(b) Find the probability that Nicola obtains a 6 for the second time on her 8th throw.

(2)
题目中文翻译

Nicola 掷一枚普通的均匀六面骰子。随机变量 XX 表示她掷出 6 所需的掷骰次数。

(a) 求 P(X<8)\mathrm{P}(X<8)。

(2)

(b) 求 Nicola 在第 8 次掷骰时第二次掷出 6 的概率。

(2)
原卷通用作答要求

Answer all questions. You should use a calculator where appropriate.

You must show all necessary working clearly; no marks will be given for unsupported answers from a calculator.

Give non-exact numerical answers correct to 3 significant figures, or 1 decimal place for angles in degrees, unless a different level of accuracy is specified in the question.

请回答所有题目。适当时应使用计算器。必须清楚展示所有必要的计算过程;仅给出计算器答案而没有过程,不予评分。除非题目另有精度要求,非精确数值答案保留 3 位有效数字,以度为单位的角度保留 1 位小数。

解答

(a)

解法一

思路

展开

小于8即前7次内出现6,取前7次都没有6的补集。

答题过程

展开

The complement is that none of the first seven throws is a 6. Hence

P(X<8)=1−(56)7=0.720918…≈0.721.\mathrm{P}(X<8)=1-\biggl(\frac56\biggr)^7 =0.720918\ldots\approx\boxed{0.721}.

解法二

思路

展开

也可以按首次掷出6在第1至第7次分类,求等比数列的和。

答题过程

展开 P(X<8)= ∑r=17(56)r−116= 161−(5/6)71−5/6= 1−(56)7≈ 0.721.\begin{align*} \mathrm{P}(X<8) =&\,\sum_{r=1}^{7}\biggl(\frac56\biggr)^{r-1}\frac16\\[4mm] =&\,\frac16\frac{1-(5/6)^7}{1-5/6}\\[4mm] =&\,1-\biggl(\frac56\biggr)^7\\[4mm] \approx&\,\boxed{0.721}. \end{align*}

(b)

解法一

思路

展开

前7次恰有一次6,第8次为6。第一次6有7个可选位置。

答题过程

展开 P= (71)16(56)616= 0.0651190…≈ 0.0651.\begin{align*} \mathrm{P}=&\,\binom71\frac16\biggl(\frac56\biggr)^6\frac16\\[4mm] =&\,0.0651190\ldots\\[4mm] \approx&\,\boxed{0.0651}. \end{align*}